# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 2.1.1 Markovian dynamics after a thermal quench
In the context of Markovian relaxation, we associate a Mpemba effect with cases where preparing a system in equilibrium at a hot temperature $T_{h}$ speeds up the cooling process towards a bath temperature $T_{b}$ in comparison to an analogous quench initiated from a colder temperature $T_{b}$ such that $T_{h}>T_{c}>T_{b}$. We refer to the temporal evolution of a system initially prepared in equilibrium at $T_{h}$ and $T_{c}$ as $p_{h}(\vec{x},t)$ and $p_{c}(\vec{x},t)$, respectively. In other words, we set $p_{\{h,c\}}(\vec{x},t=0)=\pi\left(\vec{x},T_{\{h,c\}}\right)$ and follow the out-of-equilibrium dynamics of these distributions towards $\pi(\vec{x},T_{b})$. If $p_{h}(\vec{x},t)$ converges “more quickly” to $\pi(\vec{x};T_{b})$ than $p_{c}(\vec{x},t)$, then a ME exists for this system.
@@
respectively. Assuming ergodicity (which, for rate matrices, means irreducibility of $W$), the Perron-Frobenius theorem ensures convergence to a unique steady state [68]. This implies that the dominant eigenvalue (the one with the largest real part) is identically zero and nondegenerate. It is convenient to label the eigenvalues in the order $0=\lambda_{1}>\lambda_{2}\geq\dots\geq\lambda_{N}$. The right eigenvector $\vec{v}_{1}$ represents the steady state; the DB condition then implies that $\vec{v}_{1}=\vec{\pi}(T_{b})$.
@@
representing the projection of the initial condition on the $i$-th left eigenvector. We underline that the sign of $\vec{u}_{i}$ is arbitrary, but it determines the sign of $a_{i}$. Note that the projection coefficient $a_{i}$ is in terms of the left eigenvectors $\vec{u}_{i}$, whereas the expansion of $\vec{p}$ is in terms of the right eigenvectors $\vec{v}_{i}$. If W is symmetric, then $\vec{u}_{i}=\vec{v}_{i}$; otherwise, they are different. In general, the eigenvalues are biorthogonal and can be normalized so that $\vec{u}_{i}\cdot\vec{v}_{i}=\delta_{ij}$. The eigenvector decomposition highlights how all the amplitudes of the non-dominant eigenvectors decay exponentially in time, with a rate proportional to the (negative) eigenvalues $\lambda_{i}$.
Assume for simplicity that $\lambda_{2}$ is strictly larger than $\lambda_{3}$ or, equivalently, that $|\lambda_{2}|<|\lambda_{3}|$. Then, after a long enough time, the time-dependent probability is approximately
$\vec{p}(t,T_{0},T_{b})\approx\vec{\pi}(T_{b})+a_{2}(T_{0},T_{b})e^{\lambda_{2}(T_{b})t}\vec{v}^{\,(2)}(T_{b}),$ (14)
@@
${D}[\vec{p}(t,T_{0},T_{b}),\vec{\pi}(T_{b})]\approx\left|a_{2}(T_{0},T_{b})\right|\,e^{\lambda_{2}(T_{b})t}\sum_{i}\left|v_{i}^{(2)}(T_{b})\right|.$ (15)
@@ 2.1.2 Example: Finite number of microscopic configurations
As a first example, let us demonstrate the effect in the simplest possible system, a three-state system [3]. The main advantages of the three-state system are that the distribution space, the equilibrium locus, and the relaxation trajectories in the distribution space can all be plotted; see, e.g., Fig. 5.
Returning to a three-state example, we first need to discuss the rate matrix associated with the system. A common temperature dependence of the rate matrix elements that satisfy the detailed balance condition Eq. (7) is the Arrhenius form [69],
$W_{i\!j}(T_{b})=\Gamma\,e^{-\frac{B_{i\!j}-E_{j}}{T_{b}}},$ (17)
where $\Gamma$ is an amplitude regulating the timescale of the system and $B_{i\!j}=B_{ji}$ represents a potential barrier separating the two states (hence $B_{i\!j}\geq E_{i}$ and $B_{i\!j}\geq E_{j}$). Such a form is convenient, as it connects to an underlying continuous-space model, which can be related to the discrete model resulting from a spatial coarse-graining (Fig. 4).
@@
Figure 4: (a) Sketch of an energy landscape with three metastable states. Each state’s energy is denoted by $E_{i}$, and the barrier between two states is denoted by $B_{i\!j}$. (b) An effective description of the energy landscape in (a) as a three-state system, where $R_{i\!j}$ denotes the transition rate from $j$ to $i$ and is given by Eq. 17. These figures are reprinted from [3].
@@
Figure 5: (a) The probability distribution among the three states can be described by the vector $\vec{p}=(p_{1},p_{2},p_{3})$, and all possible values of $\vec{p}$ form a simplex in $(p_{1},p_{2},p_{3})$ (shaded triangle). The curved line is the quasistatic locus, namely the set of Boltzmann distributions $\vec{\pi}(T)$ corresponding to different temperatures from 00 (blue end) to $\infty$ (red end). (b) The projection coefficient $a_{2}(T_{0},T_{b})$ at $T_{b}=0.1$ exhibits a nonmonotonic dependence on $T_{0}$ with $\partial_{T_{0}}a_{2}(T_{0},T_{b})=0$ at $T_{0}\simeq 0.42$, which is a sufficient condition to prove the existence of a Mpemba effect in this system. (c) The set of all normalized probability distributions forms a simplex — the dashed blue triangle. The red solid curve is the quasistatic locus $\vec{\pi}(T)$. The dashed arrows are the two right eigenvectors of the rate matrix, $\mathbf{W}$, $\vec{v}_{2}$, and $\vec{v}_{3}$, associated with slow (green) and fast (blue) relaxation modes. The dotted lines represent the relaxation process, $\vec{p}^{\,c}(t)$ starting at $T_{c}=0.42$ (orange) and $\vec{p}^{\,h}(t)$ starting at $T_{h}=1.3$ (purple). In this example with $T_{b}=0.1$, the ratio $\frac{a_{2}(T_{0},T_{b})\lambda_{2}(T_{b})}{a_{3}(T_{0},T_{b})\lambda_{3}(T_{b})}\lesssim 10^{-3}$ for every $T_{0}>0$, implying that in the first stages of the relaxation the out-of-equilibrium relaxation follows along $\vec{v}_{3}$. The inset shows the distances ${D}[\vec{p}^{\,h}(t),\vec{\pi}(T_{b})]$ and ${D}[\vec{p}^{\,c}(t),\vec{\pi}(T_{b})]$ both decrease with time, where the initially hot system starts at a larger distance, but after some time, its distance from equilibrium is smaller than that of the initially cold system. Panels (a) and (c) are reprinted from [3].
@@
To observe a Mpemba effect, consider a specific example where the energies of the three states are given by $E_{1}=0$, $E_{2}=0.1$ and $E_{3}=0.7$, and the barriers between them are chosen to be $B_{12}=1.5$, $B_{13}=0.8$ and $B_{23}=1.2$. The distribution space of this system is identified by the set of all probability vectors $\vec{p}=(p_{1},p_{2},p_{3})$ with components satisfying $0\leq p_{i}\leq 1$ and $p_{1}+p_{2}+p_{3}=1$. Thus, the distribution space is a two-dimensional simplex (triangle) whose vertices are at $(1,0,0)$, $(0,1,0)$ and $(0,0,1)$; see Fig.5a. The quasistatic locus $\vec{\pi}(T)$ of this system is shown as a gradient-color curve in distribution space in Fig. 5a. This parametric curve represents all the Boltzmann equilibrium configurations at temperatures $T\in[0,\infty]$. At zero temperature, the system is in its lowest-energy configuration (here, $E_{1}$), which corresponds to $\vec{\pi}(T=0)=(1,0,0)$. At infinite temperature, the Boltzmann distribution is uniform over the states, with $\vec{\pi}(T=\infty)=(\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}})$.
@@
To prove the existence of the Mpemba effect in this setup, we choose a bath temperature $T_{b}=0.1$ and analyze the $a_{2}(T_{0},T_{b}=0.1)$ coefficient during cooling quenches. In particular, if its dependence on the initial temperature $T_{0}>T_{b}$ is nonmonotonic, then $a_{2}$ reaches an extremum at some temperature, which we set to be our cold temperature; i.e., $\partial_{T_{0}}a_{2}|_{T_{0}\equiv T_{c}}=0$. This is the case for our system (Fig. 5b) with $T_{c}\simeq 0.42$ maximizing locally $|a_{2}|$. Thus, we can find a hot temperature $T_{h}>T_{c}$ such that $|a_{2}(T_{0}=T_{h})|<|a_{2}(T_{0}=T_{c})|$. This is verified in the example shown in Fig. 5 for $T_{h}=1.3$, which proves that the Mpemba effect exists in that system. In Fig. 5c, we show the trajectories of the cooling quench starting from this set of hot and cold temperatures, highlighting how the last stage of the evolution occur along a direction parallel to $\vec{v}_{2}$. Additionally, we evaluate the “entropic distance measure”
${D}[\vec{p}(t,T_{0},T_{b}),\vec{\pi}(T_{b})]=\sum_{i=1}^{M}\frac{E_{i}\left[p_{i}(t;T_{0},T_{b})-\pi_{i}(T_{b})\right]}{T_{b}}+p_{i}(t,T_{0},T_{b})\ln p_{i}(t,T_{0},T_{b})-\pi_{i}(T_{b})\ln\pi_{i}(T_{b}),$ (19)
for the two initial conditions at $T_{h}$ and $T_{c}$. In the inset, a plot of the distances from equilibrium of the two evolutions cross at $t^{*}\approx 9$, highlighting how the initially hotter system has become closer to the target Boltzmann equilibrium at $T_{b}$. More formally, ${D}[\vec{p}(t,T_{h},T_{b}),\vec{\pi}(T_{b})]>{D}[\vec{p}(t,T_{c},T_{b}),\vec{\pi}(T_{b})]$ for $t>t^{*}$, proving the existence of the Mpemba effect.
@@
One might think a two-state system would be an even simpler system to demonstrate the Mpemba effect, but as we next discuss, this is not the case – there cannot be a Mpemba effect in a Markovian system with only two states. The main reason for this fact is that for such a system, distribution space is one dimensional, since the probability vector can be written as $\vec{p}=(p,1-p)$. The rate matrix can be written as
@@
$p(t)=\frac{1}{a/b+1}+\left(p(t=0)-\frac{1}{a/b+1}\right)e^{-(a+b)t},$ (24)
@@ 2.1.3 Example: Brownian particle in a potential
Figure 6: (a) Sketch of an energy landscape in a high-dimensional configuration space as a 2D energy funnel. The funnel is cut open for illustrative purposes. The solid arrow represents a fast relaxation, and the dashed arrow a slow one. (b) Example 1D energy landscape to demonstrate the Mpemba effect in the corresponding Fokker–Planck dynamics. The left well coarse-grains to a metastable state, whereas the right well coarse-grains to the lowest-energy state. The width of the basin of attraction of the deeper well is significantly greater than that of the shallower well. (c) Decay of the distance measure $D[p(x,t),\pi(x;T_{b})]$ (see Eq. 30) from the target Boltzmann equilibrium (here $T_{b}=0.45$). A system prepared at an initially hotter temperature ($T_{h}=10$, red line) relaxes faster than one prepared at a colder temperature ($T_{c}=1.38$, blue line). (d) The Boltzmann distributions at different temperatures. After the quench, both initial distributions relax toward the final equilibrium distribution (dotted green line). The initially colder (dashed blue line) system is more populated in the lowest well than the initially hot system (solid red line); however, after a short relaxation, the hot system ends up with a higher population in the lowest well, on account of the fast relaxation from initial states within its basin. This implies that, statistically, the relaxation from an initially hot system is faster than that from an initially colder system, and the Mpemba effect occurs. Panel (a) is reproduced from [3].
@@
where $\gamma$ is the drag coefficient and the diffusion constant $D(T_{b})=k_{B}T_{b}/\gamma$ is related to the bath temperature through the Einstein-Smoluchowski relation, and $\xi(t)$ is white Gaussian noise obeying $\xi(t)\right>=0$ and $\xi(t)\xi(t^{\prime})\right>=\delta(t-t^{\prime})$.
The associated Fokker-Planck equation describes the time evolution of the probability density to find the particle at a position $x$, which we write as $p(x,t)$. Here, $p$ is a density in $x$ but not in $t$, which is simply a parameter. The time evolution obeys
$\partial_{t}p(x,t)=-\frac{1}{\gamma}\partial^{2}_{x}U(x)\partial_{x}p(x,t)+D(T_{b})\partial_{x}^{2}p(x,t).$ (26)
We can rewrite Eq. (26) as $\partial_{t}p(x,t)=\mathcal{L}_{T_{b}}p(x,t)$, introducing the Fokker-Planck operator,
$\mathcal{L}_{T_{b}}=\partial_{x}^{2}U(x)+T_{b}\partial_{x}^{2},$ (27)
where we assumed $\gamma=1$ and $k_{\rm B}=1$ for simplicity. One can easily verify that in the long-time limit, the system relaxes toward the Boltzmann distribution,
$\pi(x,T_{b})=\frac{e^{-U(x)/T_{b}}}{\mathcal{Z}(T_{b})},$ (28)
@@
In the specific example that we are considering in order to demonstrate the Mpemba effect, there is a double-well potential $U(x)$ in the interval $x\in[0,80]$, as illustrated in Fig. 6b. The potential is characterized by a local minimum at $x\approx 10$ and an absolute minimum at $x\approx 30$.
@@
where $a_{i}$ is a coefficient retaining information about the initial condition and hence depends explicitly on $T_{0}$. We set as a target bath temperature for the cooling process $T_{b}=0.45$. At this temperature, the distribution is mostly peaked around the second (deeper) well; see the green line in Fig. 6d. In the long-time limit, $a_{2}$ controls the relaxation; hence, we can study its dependence on the initial temperature $T_{0}$ and search for a nonmonotonic dependence on it. For the specific potential chosen here with $T_{b}=0.45$, the projection coefficient $|a_{2}(T_{0},T_{b})|$ has a local maximum as a function of $T_{0}$ at $1.38$. We choose the cold temperature to be $T_{c}=1.38$, as this choice maximizes the difference between $a_{2}(T_{c},T_{b})$ and $a_{2}(T_{h},T_{b})$ coefficients. (This idea is illustrated in the discussion on the strong Mpemba effect, Fig. 9a.)
@@
As in the discrete case, we can analyze the evaluation of a distance measure for a probability density function. Instead of the $L_{1}$ normed used there, here we choose to use the Kullback-Leibler divergence,
@@
In this specific choice of $U(x)$, there is a simple physical interpretation for the Mpemba effect: the potential has two wells, separated by a relatively high barrier. In addition, the deep well has a wider basin of attraction. At the lowest temperature we chose, $T_{b}=0.45$, the equilibrium distribution is tightly concentrated at the bottom of the two wells, where the ratio between the accumulated probability in the two wells is dictated by the energy difference between the energy minima of the wells, in comparison to $T_{b}$ (see green line in Fig. 6d). For the cold temperature $T_{c}$, the distribution is still peaked at the two minima, but the ratio between the probabilities within each well is different. Lastly, in the high temperature (red line in Fig. 6d), the probability is almost uniformly distributed across all space. The slowest relaxation mode in this case has a clear physical interpretation: it corresponds to diffusive motion from one well to the other. However, as the equilibrium distributions of both $T_{b}$ and $T_{h}$ have the same probability to be within each well, the relaxation takes place without the slowest mode. In contrast, the two wells are not balanced in terms of accumulated probabilities when initiating the system at $T_{c}$. Thus, in this case, a significant portion of the dynamics occurs in the slowest relaxation mode. Therefore, the relaxation from $T_{h}$ is faster than from $T_{c}$, as can be seen in Fig. 6c.
While this specific choice of $U(x)$ allows a clear understanding of the geometric mechanism behind the Mpemba effect, it is important to stress that the effect can be observed in less heavily engineered potentials. A simpler double-well potential can also exhibit the Mpemba effect [70], as seen in an experiment [6, 64]. The Mpemba effect is also predicted in an even simpler asymmetric, single-well potential [71].

### v3

@@ 2.1.1 Markovian dynamics after a thermal quench
In the context of Markovian relaxation, we associate a Mpemba effect with cases where preparing a system in equilibrium at a hot temperature $T_{h}$ speeds up the cooling process towards a bath temperature $T_{b}$ in comparison to an analogous quench initiated from a colder temperature $T_{b}$ such that $T_{h}>T_{c}>T_{b}$. We refer to the temporal evolution of a system initially prepared in equilibrium at $T_{h}$ and $T_{c}$ as $p_{h}(\vec{x},t)$ and $p_{c}(\vec{x},t)$, respectively. In other words, we set $p_{\{h,c\}}(\vec{x},t=0)=\pi\left(\vec{x},T_{\{h,c\}}\right)$ and follow the out-of-equilibrium dynamics of these distributions towards $\pi(\vec{x},T_{b})$. If $p_{h}(\vec{x},t)$ converges “more quickly” to $\pi(\vec{x};T_{b})$ than $p_{c}(\vec{x},t)$, then a Mpemba effect exists for this system.
@@
respectively. Assuming ergodicity (which, for rate matrices, means irreducibility of $W$), the Perron-Frobenius theorem ensures convergence to a unique steady state [68]. This implies that the dominant eigenvalue, the one with the largest real part, is identically zero and nondegenerate. It is convenient to label the eigenvalues in the order $0=\lambda_{1}>\lambda_{2}\geq\dots\geq\lambda_{N}$. The right eigenvector $\vec{v}_{1}$ represents the steady state; the DB condition then implies that $\vec{v}_{1}=\vec{\pi}(T_{b})$.
@@
representing the projection of the initial condition on the $i$-th left eigenvector. We underline that the sign of $\vec{u}_{i}$ is arbitrary, but it determines the sign of $a_{i}$. Note that the projection coefficient $a_{i}$ is in terms of the left eigenvectors $\vec{u}_{i}$, whereas $\vec{p}$ is expanded over its right eigenvectors $\vec{v}_{i}$. If W is symmetric, then $\vec{u}_{i}=\vec{v}_{i}$; otherwise, they are different. In general, the eigenvectors are biorthogonal and can be normalized so that $\vec{u}_{i}\cdot\vec{v}_{i}=\delta_{ij}$. The eigenvector decomposition highlights how all the amplitudes of the non-dominant eigenvectors decay exponentially in time, with a rate proportional to the (negative) eigenvalues $\lambda_{i}$.
Assume for simplicity that $\lambda_{2}$ is strictly larger than $\lambda_{3}$ or, equivalently, that $|\lambda_{2}|<|\lambda_{3}|$. Then, after a long-enough time, the time-dependent probability is approximately
$\vec{p}(t,T_{0},T_{b})\approx\vec{\pi}(T_{b})+a_{2}(T_{0},T_{b})e^{\lambda_{2}(T_{b})t}\vec{v}_{2}(T_{b}),$ (14)
@@
${D}[\vec{p}(t,T_{0},T_{b}),\vec{\pi}(T_{b})]\approx\left|a_{2}(T_{0},T_{b})\right|\,e^{\lambda_{2}(T_{b})t}\sum_{k}\left|\left(\vec{v}_{2}\right)_{k}(T_{b})\right|.$ (15)
@@ 2.1.2 Example: Finite number of microscopic configurations
Figure 4: (a) Sketch of an energy landscape with three metastable states. Each state’s energy is denoted by $E_{i}$, and the barrier between two states is denoted by $B_{i\!j}$. (b) An effective description of the energy landscape in (a) as a three-state system, where $R_{i\!j}$ denotes the transition rate from $j$ to $i$ and is given by Eq. 17. Source: Reprinted with permission from [3].
Let us first demonstrate the effect in the simplest possible system, a three-state system [3], as depicted in Fig. 4. The main advantages of the three-state system are that the distribution space, the equilibrium locus, and the relaxation trajectories in the distribution space can all be plotted; see, e.g., Fig. 5.
@@
Figure 5: (a) The probability distribution among the three states can be described by the vector $\vec{p}=(p_{1},p_{2},p_{3})$, and all possible values of $\vec{p}$ form a simplex in $(p_{1},p_{2},p_{3})$ (shaded triangle). The curved line is the quasistatic locus, namely the set of Boltzmann distributions $\vec{\pi}(T)$ corresponding to different temperatures from 00 (blue end) to $\infty$ (red end). (b) The projection coefficient $a_{2}(T_{0},T_{b})$ at $T_{b}=0.1$ exhibits a nonmonotonic dependence on $T_{0}$, with $\partial_{T_{0}}a_{2}(T_{0},T_{b})=0$ at $T_{0}\simeq 0.42$, which is a sufficient condition to prove the existence of a Mpemba effect in this system. (c) The set of all normalized probability distributions forms a simplex, the dashed blue triangle. The red solid curve is the quasistatic locus $\vec{\pi}(T)$. The dashed arrows are the two right eigenvectors of the rate matrix, $\mathbf{W}$, $\vec{v}_{2}$, and $\vec{v}_{3}$, associated with slow (green) and fast (blue) relaxation modes. The dotted lines represent the relaxation process, $\vec{p}^{\,c}(t)$ starting at $T_{c}=0.42$ (orange) and $\vec{p}^{\,h}(t)$ starting at $T_{h}=1.3$ (purple). In this example with $T_{b}=0.1$, the ratio $\frac{a_{2}(T_{0},T_{b})\lambda_{2}(T_{b})}{a_{3}(T_{0},T_{b})\lambda_{3}(T_{b})}\lesssim 10^{-3}$ for every $T_{0}>0$, implying that, in the first stages of the relaxation, the out-of-equilibrium relaxation is along $\vec{v}_{3}$. The inset shows that the distances ${D}[\vec{p}^{\,h}(t),\vec{\pi}(T_{b})]$ and ${D}[\vec{p}^{\,c}(t),\vec{\pi}(T_{b})]$ both decrease with time. The initially hot system starts at a larger distance, but after some time, its distance from equilibrium is smaller than that of the initially cold system. Source: Panels (a) and (c) are reprinted with permission from [3].
To define a three-state example, we first need to specify the rate matrix describing transitions between states. A common temperature dependence of the rate matrix elements that satisfy the detailed balance condition Eq. (7) is the Arrhenius form [69],
$W_{i\!j}(T_{b})=\Gamma\,e^{-\frac{B_{i\!j}-E_{j}}{k_{B}T_{b}}},$ (17)
where $\Gamma$ is an amplitude that sets the timescale of the system and $B_{i\!j}=B_{ji}$ represents a potential barrier that separates the two states (hence $B_{i\!j}\geq E_{i}$ and $B_{i\!j}\geq E_{j}$). Such a form is convenient, as it emerges naturally when spatially coarse graining a continuous-space model; see Fig. 4.
@@
To observe a Mpemba effect, consider a specific example where the energies of the three states are given by $E_{1}=0$, $E_{2}=0.1$, and $E_{3}=0.7$, and the barriers between them are chosen to be $B_{12}=1.5$, $B_{13}=0.8$, and $B_{23}=1.2$. The distribution space of this system is identified by the set of all probability vectors $\vec{p}=(p_{1},p_{2},p_{3})$ with components satisfying $0\leq p_{i}\leq 1$ and $p_{1}+p_{2}+p_{3}=1$. Thus, the distribution space is a two-dimensional simplex (triangle) whose vertices are at $(1,0,0)$, $(0,1,0)$ and $(0,0,1)$; see Fig.5a. The quasistatic locus $\vec{\pi}(T)$ of this system is shown as a gradient-color curve in distribution space in Fig. 5a. This parametric curve represents all the Boltzmann equilibrium configurations at temperatures $T\in[0,\infty]$. At zero temperature, the system is in its lowest-energy configuration (here, $E_{1}$), which corresponds to $\vec{\pi}(T=0)=(1,0,0)$. At infinite temperature, the Boltzmann distribution is uniform over the states, with $\vec{\pi}(T=\infty)=(\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}})$.
@@
To prove the existence of the Mpemba effect in this setup, we choose a bath temperature $T_{b}=0.1$ and analyze the $a_{2}(T_{0},T_{b}=0.1)$ coefficient during cooling quenches. In particular, if its dependence on the initial temperature $T_{0}>T_{b}$ is nonmonotonic, then $a_{2}$ reaches an extremum at some temperature, which we set to be our cold temperature; i.e., $\partial_{T_{0}}a_{2}|_{T_{0}\equiv T_{c}}=0$. This is the case for our system (Fig. 5b) with $T_{c}\simeq 0.42$ maximizing locally $|a_{2}|$. Thus, we can find a hot temperature $T_{h}>T_{c}$ such that $|a_{2}(T_{0}=T_{h})|<|a_{2}(T_{0}=T_{c})|$. This is verified in the example shown in Fig. 5 for $T_{h}=1.3$, which proves that the Mpemba effect exists in that system. In Fig. 5c, we show the trajectories of the cooling quench starting from this set of hot and cold temperatures, highlighting how the last stage of the evolution occurs along a direction parallel to $\vec{v}_{2}$. Additionally, we evaluate the “entropic distance measure”
${D}[\vec{p}(t,T_{0},T_{b}),\vec{\pi}(T_{b})]=\sum_{i=1}^{M}\left(\frac{E_{i}}{k_{B}T_{b}}\right)\left[p_{i}(t;T_{0},T_{b})-\pi_{i}(T_{b})\right]+p_{i}(t,T_{0},T_{b})\ln p_{i}(t,T_{0},T_{b})-\pi_{i}(T_{b})\ln\pi_{i}(T_{b}),$ (19)
for the two initial conditions at $T_{h}$ and $T_{c}$. In the inset, a plot of the distances from equilibrium of the two evolutions cross at $t^{*}\approx 9$, highlights how the initially hotter system has become closer to the target Boltzmann equilibrium at $T_{b}$. More formally, ${D}[\vec{p}(t,T_{h},T_{b}),\vec{\pi}(T_{b})]>{D}[\vec{p}(t,T_{c},T_{b}),\vec{\pi}(T_{b})]$ for $t>t^{*}$, proving the existence of the Mpemba effect.
@@
One might think a two-state system would be an even simpler setting to demonstrate the Mpemba effect, but there can be no Mpemba effect in a Markovian system with only two states. To understand why, recall that the distribution space is one dimensional, since the probability vector can be written as $\vec{p}=(p,1-p)$. The rate matrix can be written as
@@
$p(t)=\frac{1}{a/b+1}+\left(p(0)-\frac{1}{a/b+1}\right)e^{-(a+b)t},$ (24)
@@ 2.1.3 Example: Brownian particle in a potential
Figure 6: (a) Sketch of an energy landscape in a high-dimensional configuration space as a 2D energy funnel, cut open for illustrative purposes. The solid arrow represents a fast relaxation, and the dashed arrow a slow one. (b) Example 1D energy landscape to demonstrate the Mpemba effect in the corresponding Fokker–Planck dynamics. The left well coarse-grains to a metastable state, whereas the right well coarse-grains to the lowest-energy state. The width of the basin of attraction of the deeper well is significantly greater than that of the shallower well. (c) Decay of the distance measure $D[p(x,t),\pi(x,T_{b})]$, as in Eq. (29), from the target Boltzmann equilibrium (here $T_{b}=0.45$). A system prepared at an initially hotter temperature ($T_{h}=10$, red line) relaxes faster than one prepared at a colder temperature ($T_{c}=1.38$, blue line). (d) The Boltzmann distributions at different temperatures. After the quench, both initial distributions relax toward the final equilibrium distribution (dotted green line). The initially colder (dashed blue line) system is more populated in the lowest well than the initially hot system (solid red line); however, after a short relaxation, the hot system ends up with a higher population in the lowest well, on account of the fast relaxation from initial states within its basin. This implies that, statistically, the relaxation from an initially hot system is faster than that from an initially colder system, and the Mpemba effect occurs. Source: Panel (a) is reprinted with permission from [3].
@@
where $\gamma$ is the drag coefficient, $D$ the diffusion constant, and $\xi(t)$ is white Gaussian noise obeying $\xi(t)\right>=0$ and $\xi(t)\xi(t^{\prime})\right>=\delta(t-t^{\prime})$. The diffusion constant, $D(T_{b})=k_{B}T_{b}/\gamma$, is related to the bath temperature through the Einstein-Smoluchowski relation.
The associated Fokker-Planck equation describes the time evolution of the probability density $p(x,t)$ to find the particle at a position $x$ at time $t$. Note that $p$ is a density in $x$ but not in $t$, which is simply a parameter. The time evolution obeys
$\partial_{t}p(x,t)=-\frac{1}{\gamma}\partial_{x}[(\partial_{x}U(x))p(x,t)]+D(T_{b})\partial_{x}^{2}p(x,t)=\mathcal{L}_{T_{b}}p(x,t),$ (26)
where $\mathcal{L}_{T_{b}}$ is the Fokker-Planck operator. In the long-time limit, the system relaxes toward the Boltzmann distribution,
$\pi(x,T_{b})=\frac{e^{-U(x)/k_{B}T_{b}}}{\mathcal{Z}(T_{b})},$ (27)
@@
Figure 6b illustrates how a Mpemba effect can arise, using a double-well potential $U(x)$ in the interval $x\in[0,80]$. The potential has a local minimum at $x\approx 10$ and an absolute minimum at $x\approx 30$.
@@
where $a_{i}$ is a coefficient retaining information about the initial condition and hence depends explicitly on $T_{0}$. We set as a target bath temperature for the cooling process $T_{b}=0.45$. At this temperature, the distribution is mostly peaked around the second (deeper) well; see the green line in Fig. 6d. In the long-time limit, $a_{2}$ controls the relaxation; hence, we can study its dependence on the initial temperature $T_{0}$ and search for a nonmonotonic dependence on it. For the specific potential chosen here with $T_{b}=0.45$, the projection coefficient $|a_{2}(T_{0},T_{b})|$ has a local maximum as a function of $T_{0}$ at $1.38$. We choose the cold temperature to be $T_{c}=1.38$, as this choice maximizes the difference between $a_{2}(T_{c},T_{b})$ and $a_{2}(T_{h},T_{b})$ coefficients. (This idea is later illustrated in the discussion of the strong Mpemba effect; see Fig. 9a.)
@@
As in the discrete case, we can analyze the distance measure between probability density functions. Instead of the $L_{1}$ norm used there, we choose here to use the Kullback-Leibler divergence,
[…]
@@
In this specific choice of $U(x)$, there is a simple physical interpretation for the Mpemba effect: the potential has two wells separated by a relatively high barrier, and the deep well has a wider basin of attraction. At the lowest temperature, $T_{b}=0.45$ and the equilibrium distribution is tightly concentrated at the bottom of the two wells. The ratio between the accumulated probability in the two wells is dictated by the energy difference between the energy minima of the wells, relative to $T_{b}$; see green line in Fig. 6d. For the cold temperature $T_{c}$, the distribution is still peaked at the two minima, but the ratio between the probabilities within each well is different. Lastly, in the high temperature (red line in Fig. 6d), the probability is almost uniformly distributed across all space. The slowest relaxation mode in this case has a clear physical interpretation: it corresponds to diffusive motion from one well to the other. However, as the equilibrium distributions of both $T_{b}$ and $T_{h}$ have the same probability to be within each well, the relaxation takes place without the slowest mode. In contrast, the two wells are not balanced in terms of accumulated probabilities when initiating the system at $T_{c}$. Thus, in this case, a significant portion of the dynamics occurs in the slowest relaxation mode. Therefore, the relaxation from $T_{h}$ is faster than from $T_{c}$, as can be seen in Fig. 6c.
While this specific choice of $U(x)$ allows a clear understanding of the geometric mechanism behind the Mpemba effect, it is important to stress that the effect can be observed in less heavily engineered potentials. A simpler double-well potential can also exhibit the Mpemba effect [70], as seen in an experiment [6, 64]. The Mpemba effect is also predicted in an even simpler asymmetric, single-well potential in a box [71].